# SIRD: A Sender-Informed, Receiver-Driven Datacenter Transport Protocol

### v4

@@ 5.3. Congestion Response
We now zoom in on how each protocol manages congestion across different levels of applied stress. Figure 5 plots maximum buffering across ToR switches as a function of achieved goodput for balanced (top), core (middle), and incast (bottom) configurations. The reported goodput is the mean across all 144 hosts and reflects the rate of message delivery to applications. ToR queuing covers both downlinks and links to aggregation switches (core).
Overall, SIRD consistently achieves high goodput while causing minimal buffering across load levels, even under high levels of stress. Through informed overcommitment, SIRD makes effective use of a limited amount of credit and achieves high goodput while reducing the need for packet buffer space by up to 20×20$\times$20 × compared to Homa’s controlled overcommitment. Further, SIRD displays a more predictable congestion response compared to dcPIM and ExpressPass. As expected, based on 4(a), DCTCP and Swift cause meaningful buffering without achieving exceptional goodput. Mean queuing is qualitatively similar (appendix Figure 12).
dcPIM and ExpressPass sometimes compare favorably to SIRD but only when the workload mean message size is not small. They both struggle in terms of achieved goodput and buffering in WKa for all three configurations (left column). In WKa, $\approx 99\%$ of messages are smaller than $1\times BDP$ and are responsible for $\approx 40\%$ of the traffic. ExpressPass’s behavior is known and discussed in (Cho et al., 2017): more credit than needed may be sent for a small message which may then compete for bandwidth with productive credit of a large message. For dcPIM, the likely reason as discussed in (John K. Ousterhout, 2022b) is that to send a small message, a sender has to preempt the transmission of a larger message which can cause the receiver of the latter to remain partially idle. SIRD’s behavior is consistent across workload message sizes thanks to slight downlink overcommitment which absorbs small discontinuities in large message transmission.
In the core configuration (Figure 5 - middle row), where the core is the bottleneck, we observe that SIRD’s reactive congestion management, despite sharing the same ECN-based mechanism as DCTCP, achieves steadier behavior because it limits the number of bytes in the network. Homa grows ToR uplink queues aggressively as it does not explicitly manage core queuing which is only limited by the aggregate overcommitment level of receivers. The same applies to dcPIM but because it does not overcommit, queuing is much lower.
Last, the incast configuration (bottom row) illustrates the relative advantage of RD schemes over DCTCP and Swift.
@@
Figure 5. Maximum ToR queuing vs. achieved goodput. Configurations: Balanced (top), Core (middle), Incast (bottom).
5.4. Message Latency
Figure 6 shows the median and $99^{th}$ percentile latency slowdown for different message size ranges across configurations at 50% applied load (except WKb results which fall between the other two and can be found in appendix Figure 11). Across each workload as a whole (rightmost bar cluster ”all”), SIRD is generally on par with Homa, and often outperforms dcPIM in terms of tail latency. For small, latency-sensitive messages $<BDP$ (100KB - groups A and B), SIRD, dcPIM, and Homa offer close to hardware latency. Both in aggregate and for small messages, DCTCP and Swift perform an order of magnitude worse at the tail because they cause meaningful buffering without having a bypass mechanism like Homa.
@@ 5.4. Message Latency
Figure 6. Message latency slowdown for the median and the $99^{th}$ percentile at 50% offered application load. Each bar group contains the messages in the following size ranges: $0\leq A<MSS\leq B<1\times BDP\leq C<8\times BDP\leq D$. Also shown is the percentage of messages that belong to each group. WKc has no sub-MSS messages. ExpressPass cannot deliver 50% load in WKa and is not shown. Legend same as in Figure 5.
@@
For messages larger than $1\times BDP$ (groups C and D), SIRD comes closest to near-optimal Homa and strongly outperforms the other protocols. SIRD achieves up to 4×4$\times$4 × lower latency than dcPIM in this size range because it does not wait for multi-RTT handshakes before sending a message as discussed in §3. Note that 6(b) is comparable to Figure 3d in (Cai et al., 2022). Similarly, SIRD outperforms ExpressPass on latency because it does not take multiple RTTs to capture the full link bandwidth and it implements SRPT.
@@
The differences in latency for protocols other than dcPIM and ExpressPass, which wait before ramping up transmission, can be explained through their effective scheduling policies. That is, assuming that a protocol manages to deliver enough messages to achieve 50% goodput, lowering latency is a matter of appropriately ordering message transmission (when there are no conflicting application concerns). For example, perhaps surprisingly, DCTCP achieves equivalent or better latency than Swift in all but the incast scenarios. We argue that this is because Swift is better at fairly sharing bandwidth between messages thanks to its tighter control loop which converges faster. Latency-wise, fair sharing is inferior to FIFO or SRPT since it delays the completion of individual messages in favor of equitable progress. Along the same lines, SIRD cannot always match the latency of Homa because it approximates SRPT less faithfully as: (1) unlike Homa, a portion of the sender uplink is fair-shared (50% in this case). Note that Homa’s Linux implementation does the same to avoid starvation (Ousterhout, 2021). (2) To avoid credit accumulation in congested senders, informed overcommitment adjusts per-sender credit bucket sizes equitably.
@@
In summary, SIRD delivers messages with low latency, outperforms dcPIM, ExpressPass, DCTCP, and Swift, and, without relying on Ethernet priorities, is competitive with Homa, which nearly optimally approximates SRPT (Montazeri et al., 2018).
5.5. Sensitivity Analysis
@@
In this section, we explore SIRD’s sensitivity to its key parameters: $B$, $SThr$, and $UnschT$, as well as to the availability of Ethernet priorities. SIRD’s sensitivity to NThr is the same as DCTCP’s to $K$ (Alizadeh et al., 2010).
@@
Informed overcommitment: Equation 2 introduced the key relationship —for a simple analytical scenario— between $B$ and $SThr$, the two key parameters of informed overcommitment. Figure 7 shows the relationship in practice for Balanced WKc in terms of achieved goodput.
@@
Figure 7. Maximum goodput across values of $B$ and $SThr$.
Unscheduled transmissions: SIRD allows the unscheduled transmission of the $BDP$ prefix of messages $\leq UnschT$. Figure 8 (left) explores the sensitivity to this parameter. Setting $UnschT<BDP$ meaningfully increases median and tail latency for small messages while higher values offer no appreciable net benefit. The experiments indicated that increasing $UnschT>>BDP$ has little impact on max and mean ToR queuing, which range within $[675,750]$KB and $[365,380]$KB respectively.
The default value of $UnschT=1\times BDP$ is a satisfying compromise as large values do not yield latency benefits, yet can unnecessarily expose the fabric to coordinated traffic bursts (e.g., incast). We confirm this by running the same workload under the incast configuration which has a high concentration of $5\times BDP$ message bursts. We observe the following performance degradation when comparing $UnschT=4\times BDP$ to $UnschT=16\times BDP$: Overall $99^{th}$ percentile slowdown increases by 34% while maximum and mean ToR queuing increases by 5.7×5.7$\times$5.7 × and 1.8×1.8$\times$1.8 × respectively. These results justify our decision to introduce a size threshold above which messages are entirely scheduled.
Use of Ethernet priorities: SIRD may use a second 802.1p priority for control packets and/or unscheduled prefixes. Figure 8 (right) shows how the use of priorities impacts latency in terms of (i) possible degradation when no Ethernet priorities are available, and (ii) the benefit of prioritizing unscheduled prefixes. We observe that the slowdown of the workload as a whole is unaffected by priorities for control packets and benefits slightly by using high priority for unscheduled prefixes. Across message size groups, median latency is also largely unaffected, even for small messages, thanks to the absence of persistent queuing. Only the tail slowdown of small messages benefits from a high-priority data lane. We also observed that priority transmission of control packets increases goodput by $\approx 2\%$ as credit is delivered more predictably. Overall, the latency and utilization profiles of SIRD are not materially sensitive to the availability of Ethernet priorities. This differentiates SIRD from Homa, which fundamentally relies on priorities to bypass long queues,and dcPIM, which strongly benefits from tight control packet delivery due to its semi-synchronous nature.
@@
Figure 8. Slowdown at 50% load; Left: as a function of UnschT for WKc in the balanced configuration. Right: as a function of priority use for WKc in the balanced configuration.
6. System Implementation
We implemented SIRD as a user-level networking service on top of DPDK in 2660 lines of code and evaluated it in a real testbed. While the maturity of our testbed —10GbE NICs and a single ToR— limits the scope of the experiments, we focus on system-specific aspects and validate that (1) SIRD can have an implementation that is both simple and efficient, (2) SIRD behaves as expected under incast, provided that credits are appropriately paced (see §4.5), and (3) informed overcommitment is practical and successfully redistributes credit from congested to available senders. We do not use Ethernet priorities in the following experiments.
Incast: Given our server availability, we construct a 6-to-1 incast experiment in which 6 senders send 1MB messages to a single receiver in a closed-loop manner. The receiver responds with an 8-byte reply. A 7th sender measures queuing at the receiver’s downlink by sending 8-byte packets, waiting for the reply, and measuring the round-trip latency.
[…]
Congested sender: We evaluate informed overcommitment by designing an experiment with a congested sender. The sender starts by sending large messages to each of the four receivers. Because it can only saturate 25% of each receiver downlink, it constitutes the experiment’s bottleneck. Later, four otherwise idle hosts each send one large message to one of the four receivers (one-to-one).
Figure 10 shows the aggregate throughput at the four receiver downlinks which is at best $4\times 10=40Gbps$. When informed overcommitment is disabled ($SThr=inf$), each receiver, not knowing the state of the bottlenecked sender, gives it credit worth $1\times BDP$ out of the $B=1.5\times BDP$ it has available. This leaves each receiver with just $0.5\times BDP$ of credit to allocate to the other sender, which only yields $5Gbps$. Combined, the four receivers get $4\times(2.5+5)=30Gbps$. When SThr is properly configured ($SThr=6250=BDP/2$), the four 10Gbps links are fully utilized because the congested sender only receives the amount of credit it can use.
Figure 9. Latency CDF under 6-to-1 incast.
7. Related Work
@@
The topic of datacenter congestion control has been extensively explored due to the increasing IO speeds, the $\upmu{}$s-scale latencies, emerging programmable hardware, RPC workloads, and host-centric concerns (Alizadeh et al., 2010, 2012, 2013; Liu et al., 2021; Kumar et al., 2020; Li et al., 2019; Mittal et al., 2015; Perry et al., 2014; Le et al., 2021; Montazeri et al., 2018; Gao et al., 2015; Bai et al., 2015; Handley et al., 2017; Cho et al., 2017; Marty et al., 2019; Kalia et al., 2019; Kogias et al., 2019; Zhu et al., 2015; Addanki et al., 2022; Sun et al., 2019; Vamanan et al., 2012; Agarwal et al., 2023; Cho et al., 2020; Noormohammadpour and Raghavendra, 2018; Hong et al., 2012; Wilson et al., 2011; Lim et al., 2021; Zats et al., 2015; Cheng et al., 2014; Wang et al., 2023; Le et al., 2023; Arslan et al., 2023; Goyal et al., 2022).
@@
SD schemes like DCQCN (Zhu et al., 2015), Timely (Mittal et al., 2015), HPCC (Li et al., 2019), Bolt (Arslan et al., 2023), etc. use network signals in the form of ECN (Alizadeh et al., 2010; Zhu et al., 2015; Vamanan et al., 2012), delay (Mittal et al., 2015; Kumar et al., 2020), telemetry (Li et al., 2019; Wang et al., 2023; Zhou et al., 2021), and packet drops (Alizadeh et al., 2013) to adjust sending windows (Alizadeh et al., 2010; Kumar et al., 2020; Sun et al., 2019; Li et al., 2019) or rates (Mittal et al., 2015; Hong et al., 2012; Wilson et al., 2011). Some operate end-to-end with no or trivial switch support (Mittal et al., 2015; Kumar et al., 2020; Alizadeh et al., 2010) while others leverage novel but non-universal switch features to improve feedback quality (Li et al., 2019; Wang et al., 2023; Zhou et al., 2021) or to tighten reaction times (Arslan et al., 2023; Zats et al., 2015; Cheng et al., 2014). SIRD targets existing datacenter networks, built out of commodity switches.
@@
RD protocols attempt to avoid congestion altogether by proactively scheduling packet transmissions (Hu et al., 2020; Handley et al., 2017; Le et al., 2021; Montazeri et al., 2018; Cai et al., 2022; Cho et al., 2017; Perry et al., 2014; Gao et al., 2015; Lim et al., 2023) have already been discussed in detail. Orthogonally to the concerns of this paper, FlexPass (Lim et al., 2023) enables fair bandwidth sharing between proactive and reactive protocols. EQDS (Olteanu et al., 2022) enables the coexistence of traditional protocols like DCTCP and RDMA by moving all queuing interactions to end-hosts and using an RD approach to admit packets into the network fabric.
Figure 10. Receiver throughput under sender congestion.
@@
Previous work has followed different approaches regarding the desired flow/message scheduling policy. The traditional approach is to aim for fair bandwidth allocation at the flow level (Alizadeh et al., 2010; Kumar et al., 2020; Mittal et al., 2015; Zhu et al., 2015). Alternatively, D2TCP, D3, and Karuna (Chen et al., 2016) prioritize flows based on deadlines communicated by higher-level services. A recent line of work (Montazeri et al., 2018; Gao et al., 2015; Alizadeh et al., 2013; Bai et al., 2015; Hu et al., 2020), which includes both proactive and reactive protocols, attempts to approximate SRPT or SJF (shorted job first) scheduling in the fabric which minimizes average latency (Alizadeh et al., 2013; Bar-Noy et al., 2000). SIRD can approximate a variety of policies at receivers and senders and avoids the need to enforce fabric policies as it keeps in-network queuing minimal. We use SRPT for our evaluation because it minimizes latency in the absence of service-level information.
8. Conclusion

### v5

@@ 6.2.2 Congestion Response
We now zoom in on how each protocol manages congestion across various levels of applied stress. Figure 6 plots maximum buffering across ToR switches as a function of achieved goodput for balanced (top), core (middle), and incast (bottom) configurations. The reported goodput is the mean across all 144 hosts and reflects the rate of message delivery to applications. ToR queuing covers both downlinks and links to aggregation switches (core).
@@
Figure 6: Maximum ToR queuing vs. achieved goodput. Configurations: Balanced (top), Core (middle), Incast (bottom).
Overall, SIRD consistently achieves high goodput while causing minimal buffering across load levels, even under high levels of stress. Through informed overcommitment, SIRD makes effective use of a limited amount of credit and achieves high goodput while reducing the need for packet buffer space by up to 20×20$\times$20 × compared to Homa’s controlled overcommitment. Further, SIRD displays a more predictable congestion response compared to dcPIM and ExpressPass. As expected, based on 5(a), DCTCP and Swift cause meaningful buffering without achieving exceptional goodput. Mean queuing is qualitatively similar (appendix Figure 13).
dcPIM and ExpressPass sometimes compare favorably to SIRD but only when the mean message size is not small. They both struggle in terms of achieved goodput and buffering in WKa for all three configurations (left column). In WKa, $\approx 99\%$ of messages are smaller than $1\times BDP$ and are responsible for $\approx 40\%$ of the traffic. ExpressPass’s behavior is known and discussed in [20]: more credit than needed may be sent for a small message which may then compete for bandwidth with productive credit of a large message. For dcPIM, the likely reason as discussed in [41] is that to send a small message, a sender has to preempt the transmission of a larger message which can cause the receiver of the latter to remain partially idle. SIRD’s behavior is consistent across workload message sizes thanks to slight downlink overcommitment which absorbs discontinuities in large message transmission.
In the core configuration (Figure 6 - middle row), where the core is the bottleneck, we observe that SIRD’s reactive congestion management, despite sharing the same ECN-based mechanism as DCTCP, achieves steadier behavior because it limits the number of bytes in the network. Homa, though lacking an explicit mechanism to regulate queing at the core, does constrain peak queuing below 19MB (6MB/Tbps of switch BW) because it limits outstanding bytes to the aggregate overcommitment level of receivers. The same applies to dcPIM though queuing is much lower as it does not overcommit.
Last, the incast configuration (bottom row) illustrates the relative advantage of RD schemes over DCTCP and Swift. Homa’s results would be different if the incast optimization described in [56] was implemented and the workload consisted of request-response pairs instead of one-way messages.
@@
Figure 7: Median and $99^{th}$ percentile slowdown at 50% offered application load. Each bar group contains the messages in the following size ranges: $0\leq A<MSS\leq B<1\times BDP\leq C<8\times BDP\leq D$. Also shown is the percentage of messages that belong to each group. WKc has no sub-MSS messages. Protocols that cannot deliver 50% load are not shown.
6.2.3 Message Latency
Figure 7 shows the median and $99^{th}$ percentile slowdown for different message size ranges across configurations at 50% applied load (except WKb results which fall between the other two and can be found in appendix Figure 12). Across each workload as a whole (rightmost bar cluster "all"), SIRD is generally on par with Homa, and generally outperforms dcPIM in terms of tail latency. For small, latency-sensitive messages $<BDP$ (100KB - groups A and B), SIRD, dcPIM, and Homa offer close to hardware latency. Both in aggregate and for small messages, DCTCP and Swift perform an order of magnitude worse at the tail because they cause meaningful buffering without having a bypass mechanism like Homa.
@@ 6.2.3 Message Latency
For messages larger than $1\times BDP$ (groups C and D), SIRD comes closest to near-optimal Homa and strongly outperforms the other protocols. SIRD achieves up to 4×4$\times$4 × lower latency than dcPIM in this size range because it does not wait for multi-RTT handshakes before sending a message as discussed in §3. Note that 7(b) is comparable to Figure 3d in [16]. Similarly, SIRD outperforms ExpressPass on latency because it does not take multiple RTTs to capture the full link bandwidth, and because it implements SRPT.
@@
The differences in latency for protocols other than dcPIM and ExpressPass, which wait before ramping up transmission, can be explained through their effective scheduling policies. That is, assuming that a protocol manages to deliver enough messages to achieve 50% goodput, lowering latency is a matter of appropriately ordering the transmission of messages (assuming no conflicting application concerns). For example, DCTCP achieves equivalent or better latency than Swift in all but the incast scenarios. We argue that this is because Swift is better at fairly sharing bandwidth between messages thanks to its faster-converging control loop. Latency-wise, fair sharing is inferior to SRPT since it delays individual messages in favor of equitable progress. Along the same lines, SIRD cannot always match the latency of Homa because it approximates SRPT less faithfully: (1) unlike Homa, a portion of the sender uplink is fair-shared (50% in this case). (2) To avoid credit accumulation in congested senders, informed overcommitment adjusts per-sender credit bucket sizes equitably. This mostly impacts the latency of group C, as the bucket size may be small due to sender congestion and not grow during the lifetime of a message smaller than $8\times BDP$. Consequently, compared to near-optimal Homa in the Balanced configuration, SIRD’s $99^{th}$ percentile slowdown in group C is 1.85×1.85$\times$1.85 × and 2.68×2.68$\times$2.68 × higher at 50% and 70% application load respectively, while it outperforms the other protocols.
((a)) WKa Balanced
((b)) WKc Balanced
Figure 8: Median and $99^{th}$ percentile slowdown at 70% offered application load. Protocols that cannot deliver 70% load are not shown. Legend same as in Figure 7.
Figure 8 shows slowdown under a higher applied application load of 70% for the protocols that can deliver it. At this load level, message scheduling becomes more important and Homa’s near-optimal SRPT implementation sees its relative performance improve in some cases.
@@
In summary, SIRD delivers messages with low latency, outperforms dcPIM, ExpressPass, DCTCP, and Swift, and is competitive with Homa, which nearly optimally approximates SRPT [56].
6.2.4 Sensitivity Analysis
@@
In this section, we explore SIRD’s sensitivity to its key parameters: $B$, $SThr$, and $UnschT$, as well as to the availability of switch priority queues. SIRD’s sensitivity to $NThr$ is the same as DCTCP’s to $K$ [8].
@@
Informed overcommitment: Equation 2 introduced the linear steady-state relationship between $B$ and $SThr$, the two key parameters of informed overcommitment. Figure 9 (left) shows the measured maximum achieved goodput in Balanced WKc as a function of $B$ and $SThr$.
Figure 9: Left: Max goodput across values of $B$ and $SThr$. Right: Credit location at max goodput as a function of $SThr$. $B=1.5\times BDP$. "In-flight" is CREDIT and DATA. Credit at receivers is low at this load.
@@
With the mechanism enabled, the curves asymptotically converge to the same maximum goodput of 90Gbps. Reaching the plateau with a lower value of $B$ also demands lowering $SThr$ to reduce credit stranded at senders (Equation 2). Queuing increases with $B$ as in Figure 2 and remains stable when varying $SThr$. We selected $B=1.5\times BDP$ rather than $B=2.0\times BDP$ for our experiments because it halves the maximum buffering caused by scheduled packets from 1×1$\times$1 × to $0.5\times BDP$ ($=B-BDP$). We do not configure $SThr$ lower than $0.5\times BDP$ as we deem it unrealistic to implement in software, as it may cause unwanted marking due to batch credit arrivals. The minimum correct value for this topology is $0.15\times BDP$, as per the guidelines [8].
@@
Figure 10: Slowdown as a function of UnschT for WKa (left) and WKc (right) at 50% application load in the balanced configuration;
Unscheduled transmissions: SIRD allows the unscheduled transmission of the $BDP$ prefix of messages $\leq UnschT$. Figure 10 explores the sensitivity to this parameter using workloads WKa and WKc. Setting $UnschT=MSS$ meaningfully increases median and tail latency for messages with sizes in $[MSS,BDP]$ while higher values offer no appreciable net benefit. This also holds at 70% load. Increasing $UnschT>>BDP$ in WKc has little impact on max and mean ToR buffering, which range within $[696,747]$KB and $[363,378]$KB respectively. In contrast, because $99\%$ of messages in WKa are unscheduled, max and mean ToR buffering scale from $501$KB to $1016$KB and $220$KB to $435$KB, respectively, as UnschT increases. It is worth noting that saturating the fabric’s bandwidth, which is when queuing is maximized, is far less likely with high-request-rate workloads like WKa.
The default value of $UnschT=1\times BDP$ is a satisfying compromise as large values do not yield latency benefits yet can unnecessarily expose the fabric to coordinated traffic bursts (e.g., incast). We confirm this by running WKc under the incast configuration which has a high concentration of $5\times BDP$ message bursts. We observe the following performance degradation when comparing $UnschT=4\times BDP$ to $UnschT=16\times BDP$: Overall $99^{th}$ percentile slowdown increases by 34% while maximum and mean ToR queuing increases by 5.7×5.7$\times$5.7 × and $80\%$ respectively. These results justify our decision to introduce a size threshold above which messages are entirely scheduled.
In deployments with severe incast consisting of messages smaller than $BDP$, UnschT can be configured to be smaller than $MSS$. If, for example, 100 hosts concurrently send a $0.5\times BDP$ message to a single receiver, the latter will effectively receive $100$ small packets requesting credit. The receiver will immediately start issuing credit based on its policy and ramp-up to full bandwidth utilization one RTT after the senders transmitted. In TCP terms, this specific scenario is roughly equivalent to each sender starting with a window of one MSS, and ramping to the correct window value just one RTT later.
Use of switch priority queues: SIRD may use a second 802.1p priority level for control packets and/or unscheduled DATA. Figure 11 shows how the use of priorities impacts message slowdown for WKa and WKc in terms of (i) possible degradation when switch priority queues are unavailable, and (ii) the benefit of prioritizing unscheduled DATA.
We observe that, across message sizes, median slowdown is largely unaffected and tail slowdown benefits from high priority transmission in some cases. Small messages benefit because they bypass the small queues SIRD forms ($\approx 1\times BDP$ per link at the $99^{th}$ percentile and $\approx 0.1\times BDP$ on average). This is more evident in WKa where 99% of messages are unscheduled. Sensitivity to priorities is similar at 70% load (not shown), where, compared to 50% load, group A’s tail slowdown benefits 5% more from DATA prioritization and group B’s 6% less.
Maximum goodput increases by $1\%$ and $2.4\%$ with control packet prioritization for WKa and WKc respectively, as credit is delivered slightly more predictably. Queuing in the network is also insensitive to the use of priorities. Given SIRD’s low sensitivity to the availability of switch priority queues, we argue that it can be deployed without this dependency with little sacrifice. This differentiates SIRD from Homa, which is designed around the use of priorities to bypass long in-network queues, and dcPIM, which strongly benefits from tight control packet delivery due to its semi-synchronous nature.
Figure 11: Slowdown as a function of priority use for WKa (left) and WKc (right) at 50% application load in the balanced configuration;
7 Related Work
@@
The topic of datacenter congestion control has been extensively explored due to the increasing IO speeds, the $\upmu{}$s-scale latencies, emerging programmable hardware, RPC workloads, and host-centric concerns [8, 9, 10, 50, 44, 47, 54, 61, 46, 56, 31, 13, 36, 67, 5, 20, 51, 42, 43, 74, 4, 66, 68, 6, 22, 57, 38, 70, 48, 71, 19, 69, 45, 12, 34].
@@
SD schemes like DCQCN [74], Timely [54], HPCC [47], Bolt [12], and PTT [67] use network signals in the form of ECN [8, 74, 68], delay [54, 44], in-band telemetry [47, 69, 73], and packet drops [10] to adjust sending windows [8, 44, 66, 47] or rates [54, 38, 70]. Some operate end-to-end with no or trivial switch support [54, 44, 8, 67] while others leverage novel but non-universal switch features to improve feedback quality [47, 69, 73] or to tighten reaction times [12, 71, 19].
@@
RD protocols attempt to avoid congestion altogether by proactively scheduling packet transmissions [39, 36, 46, 56, 16, 20, 61, 31, 49, 5] and have already been discussed (§2). Orthogonally to the scope of this paper, FlexPass [49] enables fair bandwidth sharing among proactive and reactive protocols. EQDS [59] enables the coexistence of traditional protocols like DCTCP and RDMA by moving all queuing interactions to end-hosts and using an RD approach to admit packets into the network fabric.
@@
Previous work has followed different approaches regarding the desired flow/message scheduling policy. The traditional approach is to aim for fair bandwidth allocation at the flow level [8, 44, 54, 74]. Alternatively, D2TCP, D3, and Karuna [18] prioritize flows based on deadlines communicated by higher-level services. A recent line of work [56, 31, 10, 13, 39], which includes both proactive and reactive protocols, attempts to approximate SRPT or SJF (shorted job first) scheduling in the fabric which minimizes average latency [10, 14]. SIRD can approximate a variety of policies at receivers and senders and avoids the need for fabric policies as it keeps in-network queuing minimal.
8 Conclusion